# Benchmarking MedMNIST dataset on real quantum hardware

### v1

@@ 4 Conclusion
This work demonstrates the potential of QML in the context of medical imaging by benchmarking the MedMNIST datasets on real quantum hardware. We effectively trained and validated noise-resilient quantum circuits and leveraging advanced error suppression and mitigation techniques to tackle the hardware noise. The experimental results highlight the feasibility of applying QML to complex real-world datasets such as MedMNIST. In the current NISQ era, we are able to achieve promising results in terms of classification accuracy and AUC values for binary as well as multi-class datasets with very limited number of input embedding features. This study serves as an important benchmark, demonstrating the readiness of quantum computing for practical applications in healthcare. It also underscores the need for further research to enhance scalability, mitigate hardware limitations, and explore novel quantum algorithms tailored to medical data. As quantum technology continues to evolve, its integration into medical imaging workflows has the potential to transform healthcare analytics, paving the way for more advanced and efficient diagnostic tools.

### v2

@@ 4 Conclusion
This work demonstrates the potential of QML in the context of medical imaging by benchmarking the MedMNIST datasets on real quantum hardware. We effectively trained and validated noise-resilient quantum circuits using advanced error suppression and mitigation techniques to tackle hardware noise. The experimental results highlight the feasibility of applying QML to complex real-world datasets such as MedMNIST. In the current NISQ era, we can achieve promising results in terms of classification accuracy and AUC values for binary as well as multi-class datasets with a very limited number of input embedding features. For the 2-class PneumoniaMNIST data set, our best-performing QML model provides accuracy and AUC of 85.4% and 82.2%, respectively. Similarly, we obtained comparable results with the classical benchmark model in the case of the 5-class RetinaMNIST data set with accuracy and AUC of 50.5% and 66.4%, respectively. Even for the 7-class DermaMNIST dataset, our QML model achieves an accuracy of 67.7% and AUC of 69.8% showing the capability of QML for medical imaging classification. Importantly, we achieve quantum advantage both in classification performance and computational complexity compared to classical ML baselines when evaluated on the RetinaMNIST dataset using the same number of input features. This study serves as an important benchmark, demonstrating the readiness of quantum computing for practical applications in healthcare. It also underscores the need for further research to enhance scalability, mitigate hardware limitations, and explore novel quantum algorithms. As quantum technology continues to evolve, its integration into medical imaging workflows has the potential to transform healthcare analytics, paving the way for more advanced and efficient diagnostic tools.
Data availability
The MedMNIST datasets used in this benchmarking study can be accessed using https://medmnist.com/
Code availability
All the related codes are available at https://github.com/gurinder-hub/QML_MedMNIST
@@ Acknowledgments
Author contributions
G. Singh and H. Jin proposed and design the study of benchmarking MedMNIST datasets using quantum machine learning on real IBM Cleveland hardware. G. Singh and H. Jin performed the experiments under the supervision of K.M. Merz. G. Singh wrote the manuscript, with inputs and contributions from all authors.
Competing interests
The authors declare no competing interests.